OpenStudy (anonymous):

change the order of integration in the integral I=∫0 to a ∫0 to b/a sqrt(a^2-x^2) (xy)dydx and evaluate it

OpenStudy (anonymous):

is it: \[\int\limits_o^a\int\limits_0^{\frac{b}{a}\sqrt{a^2-x^2}}xy dy dx\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ya..

OpenStudy (anonymous):

y=b/a sqrt(a^2-x^2) is an elipse

OpenStudy (anonymous):